OpenStudy (kofi):

The driver of a truck slams on the brakes when he sees a tree blocking the road. The truck slows down uniformly with acceleration −6.1 m/s2 for 4.20 s, making skid marks 62.5 m long that end at the tree. With what speed does the truck then strike the tree

OpenStudy (mrnood):

you have 2 unknowns - the initial speed, u, and the final speed v so you need 2 equations to solve it use s= ut +0.5at^2 to sulve for u then use v^2= u^2 +2as to solve for v